| 
  • If you are citizen of an European Union member nation, you may not use this service unless you are at least 16 years old.

  • You already know Dokkio is an AI-powered assistant to organize & manage your digital files & messages. Very soon, Dokkio will support Outlook as well as One Drive. Check it out today!

View
 

AccessReports

Page history last edited by ACPL Staff 14 years, 10 months ago

Access (Circ and TSS) Reports:

 

May 2009

  • Current hours:

    • BBW and SS-40
    • RA-35
    • PH-23
    • NM-22
    • DS, HM, LP, JA, EP2-20
    • EP and KF-15
    • TP-7 
  • Stephen is leaving to accept a position back east.
  • Erron is returning to Idaho.
    • Remaining desk staff are working on a plan to divide his hours and assume his responsibilities.
  • Hollis and Katie took vacation breaks.

    • Ada is providing some "substitute" help.

  • Staff are working on adjusting to the Sirsi "J Client".

  • Staff are collaborating to assume responsibilities in Stephen's absence.

 

 

April 2009

  • Current hours:

    • BBW and SS-40
    • RA-35
    • PH-23
    • NM-22
    • DS, HM, LP, JA, EP2-20
    • EP and KF-15
    • TP-7
  • All staff have passed the Circ tests.

  • We "completed", at least for now, the process of streamlining holds and ILLs to use Sirsi more effectively.

    • All holds are in one place.

    • Reports are being used to notify patrons by email and generate lists of patrons to call.

    • Expiration of, and pulling of expired, holds, has been automated by Sirsi report.

  • We've stopped using the encoded "Cutter" system and began using the unencoded author's name to classify, label and sort new items.

 

 

March 2009

  • Current hours:

    • BBW and SS-40
    • RA-35
    • PH-23
    • NM-22
    • DS, HM, LP, JA, EP2-20
    • EP and KF-15
    • TP-7 
  • All staff have passed the shelving tests.

  • All Gates computers are out to the patrons.

  • Busiest month statistically since July.

    • ILL continues to grow
      • This is the second month where we've sent more items to OCLC libraries than we've received from them.
  • The library has its first netbook. Kathy, Bailey and co. are experimenting with it to see how it works and what applications it could have.

  • Erron has been helping Stephen with some IT tasks.

 

February 2009

  • Current hours:

    • BBW and SS-40
    • RA-35
    • NM-24
    • LP and PH-23
    • DS-20
    • HM-20
    • JA-20
    • EP2-20
    • EP-15
    • KF-10
    • TP-7
  • We have simplified the newspaper storage, so all newspapers can be found upstairs.

 

January 2009

  • Current hours:

    • BBW and SS-40
    • RA-35
    • NM-24
    • LP and PH-23
    • DS-20
    • HM-20
    • JA-20
    • EP2-20
    • EP-15
    • KF-10
    • TP-7
  • Statistics remain fairly flat.

  • Voltech's wiring work is done.

  • Gates computers are being put together and put out.

  • The wiki was updated to the new version of PBWiki. 

 

January 2009

  • Current hours:

    • BBW and SS-40
    • RA-35
    • LP-23
    • DS-20
    • HM-20
    • KF-20
    • NM-20
    • JA-20
    • EP2-20
    • EP-15
    • TP-7
  • A fairly busy month statistically. Due to the "item used" revision noted in last month's report, it appears as though this years January circulation number is a decrease from the previous year. However, when the change in calculation methods is accounted for, there was actually a very small increase from the previous year.

  • Voltech's wiring work began January 21st.

  • We are going to have an all-time high number of Circ Assistants, in place of the Shift Coordinator and IT jobs for now, so we can bring a large group of quality people to the library, and so we can hopefully provide for later advancement opportunities.

  • Gates computers are on their way. We delayed purchase this year due to delayed reimbursement from WSL, and wanting to get the best possible deal.

  • Erron, Jodie, Nicole, and Pat have joined us at the Circ desk.

  • We have a new PC on the Information Station desk and Stephen is working on other staff IT needs for this year.

  • Rock River Branch has a new DSL internet connection.

 

 

December 2008

 

  • Current hours:

    • BBW and SS-40
    • CK-36
    • RA-35
    • MC-20
    • SC-20
    • HM-20
    • KF-20
    • AP-20
    • EP-15
    • TP-7
  • A generally quiet month statistically.

    • The days after Xmas were super busy.
  • We have revised our stats formula for "item marked used"

    • Due to a change in how Director's Station computes stats, it was being counted twice in our totals before.
  • Several new Circulation Assistants will be joining us next month.

  • Ada and Mandy provided us with valuable help over the holidays.

  • Deb and Laura have joined us at the Circ desk. Deb is still managing the Centennial branch as well.

  • Hollis handled all the holiday book drops. Thanks!

  • Cataloging is picking up again, with a little ordering and a lot of gifts.

 

 

November 2008

 

  • Current hours:

    • BBW and SS-40
    • CK-36
    • RA-35
    • MC-20
    • SC-20
    • HM-20
    • KF-20
    • AP-20
    • EP-15
    • TP-7
  • A quiet month statistically.

  • We're very sad that Cheryl is leaving us for another position. She will be missed!

  • Amy is getting promoted to Youth Services Librarian. Congratulations!

  • Sandy's going to be doing her teaching placement in Cheyenne. We'll miss her and hope she'll be back when it's done!

  • We have completed a truth-in-labeling project to get all the shelving signs and labels to be accurate around the library. We are going to work to maintain this going forward.

  • Aaron has been a big help to us, fixing things like the end-pieces in the stacks and a drawer of our music CD shelving, and trying to improve the function of the door and gate system.

  • Hollis, Amy, Katie, Cheryl and Marianne passed the Circ tests.

  • Stephen is getting ready to buy for the final Gates Grant, which is this year.

    • The Gates foundation will no longer be funding public library computers, so we'll need to find another source for funds after this year.
  • Bernie and Stephen are continuing to work on automating the Holds/ILL process so we can keep up with the increased demand.

  • The wiring contract is finally approved and the work should begin shortly.

  • Ada continues to provide us with some very valuable substitute help.

 

 

October 2008

  • Current hours:

    • BBW and SS-40
    • CK-36
    • RA-35
    • MC-20
    • SC-20
    • HM-20
    • KF-20
    • AP-20
    • EP-15
    • TP-7
  • A slight uptick in Circ activity from last month, and a higher number than last October. This may indicate that patrons have adjusted to the new hours.

  • Cataloging numbers are back up with ordering returning somewhat.

  • The door counter was broken for most of October, so we don't have good traffic stats. It's working again as of this writing.

  • Ada is subbing for us again.

  • Sandy was the first staff member to pass the new Circ tests.

  • Bernie, Cheryl, Rekha, and Stephen visited the Laramie County Library. Their reports are here along with reports from other visits taken.

 

 

September 2008

  • Current hours:

    • BBW and SS-40
    • CK-36
    • RA-35
    • CW-21
    • MC-20
    • SC-20
    • HM-20
    • KF-20
    • AP-20
    • EP-15
    • TP-7
  • A fairly slow month by most statistical measures.

  • We have got new computers for the end of the Circ desk, and for Stephen's desk. The one on the Circ desk is our first Windows Vista PC. Stephen's PC is our first with a quad-core processor.

  • We rearranged the Circ desk with the new computer, to give Cheryl (and everyone) more room to operate, and a third phone.

  • Amy and Katie have joined us as our new Circ Assistants.

  • Stephen is still restarting the DSL connection every day to keep it going. New router installation is forthcoming.

  • We have developed a 100-question multiple-choice and 20-questions short-answer test for Circulation Assistants, and all staff are taking it. Cheryl and Stephen will be following up with staff about corrections and related training needs.

  • Evaluations and goals have been done for Bernie and Eric.

  • Cheryl has started handling missing and claimed-returned reports.

  • We are starting on a new ILL/holds workflow that will hopefully allow us to more efficiently serve patrons, and keep up with the onslaught of ILL/holds demand. We will be more fully using the automated features available in Sirsi to manage this.

 

 

August 2008

  • Current hours:

    • BBW and SS-40
    • CK-36
    • RA-35
    • CW-21
    • MC-20
    • SC-20
    • HM-20
    • Shelving subbing and misc.(MK in August)-20
    • TC-20
    • EP-15
    • TP-7
  • All Laramie staff computers are updated to Windows XP Service Pack 3.

  • With a little bit of help from Stephen over the phone, Shirley and Deb were able to install the new antivirus software on the PCs at their branches, saving the library several hundred dollars in mileage and time.

  • 2 new staff PCs are on the way, and we're working on a third.

  • Tara is moving on to Administrative Services. We're advertising to replace her position and the other vacant 20 hours. Mandy did some subbing, doing the shelving.

  • Everyone worked to keep ILL going while Bernie was on vacation. We were able to stay in the "rota" and keep sending books out to other libraries, as well as handle our own incoming items.

 

 

July 2008

  • Current hours:

    • BBW, CK and SS-40
    • RA-35
    • MC-24
    • SC-24
    • CW-20
    • HM-20
    • TC-20
    • EP-15
    • TP-7
  • Busiest overall circulation for July that I can find on record since 1987.

  • Cheryl has assumed the position of Circulation Shift Coordinator.

  • We are again without a shelver. Marianne and Sandy are filling in for now. We are going to look at how to integrate this with the Circulation Assistant tasks in the future, instead of having a separate position.

  • We now have more than 3000 music CD titles.

    • Fourth largest collection in the state, behind only Teton County, Laramie County, and Natrona County.
      • We hope to have the second largest soon.
    • We've also worked to have a high quality collection, with core works in all genres and responding to patron requests and suggestions.
  • We have switched to Trend Micro Antivirus for all PCs.

    • It provides better tracking and central monitoring of problems.
    • It's saving us some money over the alternative.
  • Rodney has ordered us a new DSL modem.

    • This should help a little with connection choppiness, hopefully.
    • We are outgrowing our connection and will need an additional one soon, however.
  • With Paul's help, we now have Nonprofit status for our Youtube account.

 

June 2008

  • Current hours:

    • BBW and SS-40
    • RA-35
    • CK-24
    • MC-20
    • SC-20
    • CW-20
    • HM-20
    • TC-20
    • EP and JK-15
    • TP-7
  • Very busy month of circulation.

  • Bernie, Cheryl and Stephen attended the WYLD meeting in Buffalo.

  • We were short staffed this month, due to making up for Chris' accrued vacation payout.

  • Chip is our new Circulation Assistant, replacing Sandy W.

  • Stephen is working to get the wiring project through the bureaucracy.

 

May 2008

  • Current hours:

BBW, CG and SS-40

RA and JS (both positions)-35

SW-29

CK-24

HM-20

TC-20

EP-15

TP-7

  • Things tend to quiet down in May, but this was busier than last year.

  • Joanna is leaving us May 24. Joni is our new shelver, and Marianne is filling her Circulation Assistant position.

  • Tara is back.

  • Sandy C. is filling Cheryl's Circulation Assistant position.

  • Sandy W. is leaving us for full-time employment elsewhere. We'll be filling her position shortly.

  • Internet service has started to be choppier lately. Better internal wiring should improve this, but we are also starting to outgrow our current bandwidth.

 

 

April 2008

  • Current hours:

BBW, CG, and SS-40

JS (Shelving plus Desk) and RA-35

SW-29

CK-24

HM-20

AG and EP-15

TP-7

  • Busiest month of circulation for the fiscal year so far.

    • Highest monthly circulation ever for music CDs, with 1673 checkouts.
  • Chris is leaving May 29. Cheryl will assume the position of Circulation Shift Coordinator one month after he leaves.

  • We are looking for one new Circulation Assistant.

  • Stephen has gotten approval to pursue the rewiring project to improve connectivity and will try to get that going ASAP.

  • The new patron server, paid for with Gates money, was installed April 18. It is running CentOS (Linux).

  • 4 new computers, paid for with Gates money, were set up at Rock River April 23. They include a replacement for the one in the library, plus 3 new public computers that will be in the public room and available 5 days a week.

  • We now have 37 training videos on Youtube:

 

 

 

March 2008

 

  • Current hours:

BBW, CG, and SS-40

JS (Shelving plus Desk) and RA-35

SW-29

CK-24

HM-20

AG and EP-15

TP-7

  • Busiest month of circulation and cataloging of the fiscal year so far.

  • We've added several new training videos to our Youtube site:

  • We are working on goals and evaluations for this year.

  • We've gotten several hundred new music CDs. We are asking for more recommendations.

  • Ada is subbing for us at the desk until Tara returns.

  • We learned at the Gates conference that there will be no more Gates hardware grants after next year, so we'll probably need to find another source for funding this.

 

February 2008

  • Current hours:

Bernie, Chris, and Stephen-40

Rekha-35

Sandy-27

Mandy-22

Hollis, Tara, Cheryl-20

Eric and Joanna-15

Tracy-7

  • Circulation and cataloging continue to be busy.

  • All the Gates equipment is out for public use, except the Rock River PCs and

the server. Rodney has set up the server and we are waiting for a good time to

install it. The Rock River PCs are still being worked on by Stephen.

  • Paul got us three older PCs from First Interstate Bank. They have replaced

the stacks OPACs and Word Processor that burned out.

  • We have found some ways to improve the disk "cloning" or "ghosting" process,

and it's working fairly well right now.

  • We have had a few bad experiences with the new Symantec Antivirus product, so

we are not doing further upgrades right now. The virus databases are still

getting updated regularly for the older product. Rodney has switched from

Symantec to Trend Micro at the Courthouse.

  • We are working on streamlining the workflow for receiving ordered items in TS.

  • Bernie is training Eric on doing packaging for ILLs.

  • Mandy is leaving us. She will be missed.

  • Mandy has passed her read-and-return and tax responsibilities on to Cheryl.

Sandy will be taking over managing the MySpace page.

  • Now that the music shelving is in place and Joanna has reorganized the CDs,

Stephen has begun this year's big music buy.

 

Comments (0)

You don't have permission to comment on this page.